# IA-TIGRIS: An Incremental and Adaptive Sampling-Based Planner for Online Informative Path Planning

### v1

@@ Abstract
Planning paths that maximize information gain for robotic platforms has wide-ranging applications and significant potential impact. To effectively adapt to real-time data collection, informative path planning must be computed online and be responsive to new observations. In this work, we present IA-TIGRIS, an incremental and adaptive sampling-based informative path planner that can be run efficiently with onboard computation. Our approach leverages past planning efforts through incremental refinement while continuously adapting to updated world beliefs. We additionally present detailed implementation and optimization insights to facilitate real-world deployment, along with an array of reward functions tailored to specific missions and behaviors. Extensive simulation results demonstrate IA-TIGRIS generates higher-quality paths compared to baseline methods. We validate our planner on two distinct hardware platforms: a hexarotor UAV and a fixed-wing UAV, each having unique motion models and configuration spaces. Our results show up to a 41% improvement in information gain compared to baseline methods, suggesting significant potential for deployment in real-world applications. Project website and video at ia-tigris.github.io.
@@ I Introduction
Robots play a vital role in gathering information from the physical world, supporting a wide range of applications such as scientific research [1, 2, 3], environmental monitoring [4, 5, 6, 7, 8], search and rescue operations [9, 10], and disaster response efforts [11, 12, 13]. By employing intelligent algorithms, robotic systems enhance the efficiency of data collection, provide valuable insights, and support well-informed decision-making processes. These autonomous robots provide unparalleled advantages in situations where human access is constrained, dangerous, or logistically challenging [14]. Moreover, the scalability and speed of autonomous information gathering robots significantly enhances the rate of information gathering by not capping the number of robots based on the number of human operators.
@@
Informative path planning (IPP) seeks to create intelligent paths for robots to execute actions that maximize their inherent potential and advantages in data gathering situations, while not violating budget constraints on the system or mission. This approach has many advantages over most coverage-based approaches which often don’t take into account the flight time of the drone or prior information about the space and don’t adapt to learned information during flight. However, solving the IPP problem is computationally expensive, being at least NP-hard even in a discrete space [15], leading to methods that only plan over a short horizon or are static during the path execution.
In this work, we introduce IA-TIGRIS (Incremental and Adaptive Tree-based Information Gathering using Informed Sampling), a sampling-based informative path planner that incrementally builds its search tree and adapts to new information online by efficiently updating and building the set of possible trajectories. This leads to global paths that are more robust to path execution disturbances, adapt to changes in the information map, and are higher quality than a single plan due to continual online refinement. Fig. 1 demonstrates IA-TIGRIS deployed on a hexarotor drone for an IPP application. Our key contributions are the following:
@@
The paper is organized as follows: Section II gives a detailed overview of related works and the contributions of this work. Section III presents the problem formulation. Our proposed approach is detailed in Section IV. The section outlines the planner’s design and its integration with system components to enable adaptive and efficient informative path planning for real-world applications. Section V provides our detailed experimental evaluation and results of the approach. The results of our field deployments are summarized in Section VI, and Section VII offers the conclusion and future work.
@@ II Related Works
One way to formulate the IPP problem is to discretize the continuous space into a fixed set of states. Many previous works have then used this discrete space to view IPP as an orienteering problem. Given a weighted-undirected graph, orienteering involves determining a Hamiltonian path over a subset of nodes with the objective of maximizing the total reward while subject to a constraint on the path cost. The orienteering problem can be thus broken down into an optimization objective that involves selecting a subset of nodes from the given graph and a traveling salesman problem, that aims to minimize the distance traveled while visiting the selected subset of nodes. This optimization objective, which dictates how the subset of nodes are selected, can be chosen based on the specific task to which orienteering is being applied. For example [16] applies it to the task of environmental monitoring in an aquatic setting while [17] applies it to the more general Constraint Satisfaction Problem (CSP).
@@
However, framing IPP as an orienteering problem leads to computationally expensive runtimes, which rapidly become intractable as the likelihood of near-optimal paths in the solution space diminishes. As the authors [18] note, the IPP problem is a NP-hard search problem. In order to overcome this, the authors develop an approximation algorithm that efficiently finds near-optimal solutions leveraging mutual information in the problem space. They do this by ensuring that the mutual information formulation is submodular in nature as discussed in [19]. This results in a diminishing returns property whereby making a new observation yields more information if fewer observations have been made up to that point, and less information if many observations have already been made. Hence, the IPP problem can be reformulated as maximizing a submodular objective function while constrained to a budget. [20] studies this formulation by developing a recursive-greedy algorithm that they show has strong theoretical guarantees.
@@
However, these mentioned approaches have only been used in applications that have relatively small search spaces such as environmental monitoring [16], water quality monitoring and analysis [21], or detection of chemical or biological plumes [22] that are then discretized into a finite set of sensing locations. As one can imagine, this becomes quickly intractable in applications that have much larger search spaces, are higher dimensional, or both. Examples of this type of situation include searching for targets of interest over a large search area, such as a missing hiker or tracking multiple dynamic targets of interest over a larger search area, such as ships at sea. In order to combat this, the authors in [16] propose using heuristics while [16, 23, 20] also suggest using greedy approaches to approximate solutions. Another method involves Mixed Integer Programming (MIP) based solutions such as [24]. But these solutions fail to capture the relationship between nodes which would not be ideal for many problem statements. The limitations in capturing the relationship in rewards among nodes along with the inefficient run times necessitate developing better exploration algorithms that would allow for robust information gathering by autonomous systems in complex problem settings.
@@
Some previous approaches have tried using receding horizon-based methods to solve the IPP problem. Local receding-horizon based solutions optimize the information-objective over a small lookahead or horizon. They rely on replanning to approach a globally optimal solution. One such approach involves a receding horizon algorithm [25] that satisfies a temporal logic specification for problem statements where safety and reliability are paramount. Another approach by [26] involves using a receding horizon planner to perform informative planning in a latent environment modeled using Gaussian processes. According to [27], receding horizon-based path planners are susceptible to becoming trapped in local optima. To address this issue, the study explores the impact of adjusting the horizon length and considers adapting it based on the remaining information. They propose a J-Horizon algorithm, which incorporates a lookahead step size to improve convergence toward better optima. Despite these improvements, receding horizon methods primarily plan optimal paths reactively and fail to fully utilize the available prior information about the environment [28].
@@
Notably, [29] compares lawnmower paths with other planning algorithms and observes that their performance is only slightly inferior to that of adaptive algorithms. Furthermore, the study concludes that graph-based search algorithms under perform compared to lawnmower patterns due to their limited ability to adapt to uncertainties in prior information. The ability to allow for revisits is very important for finding globally optimal paths, especially when the information objective is submodular or when the task requires revisits [30]. Another approach which has shown promise involves branch and bound techniques which prune suboptimal branches early in the tree search [31, 32]. However, efficiently calculating tight bounds in problems with unknown environments and high-dimensional state space is nontrivial. Previous works have solved this using Gaussian processes (GP). However, implementing GPs is application specific and a specific approach used in one work may not be transferable to another application. Secondly, the computational complexity of inference in GPs generally scale cubically with the number of observations [33]. Although methods exist to mitigate this complexity, long-horizon planning remains computationally demanding unless stationary assumptions are made about the covariance function. This is further compounded when planning in higher-dimensional state spaces.
@@
Numerous works have worked to address these drawbacks by formulating the task of information gathering in continuous space and introducing adaptive sampling-based algorithms. Examples of some of these works include [34], [35], [36, 37, 38, 39, 40, 41]. These methods employ sampling-based approaches, which involves the selection and inclusion of new states in continuous space into a tree structure. As the tree grows, the optimal path—maximizing information gain while respecting budget constraints—is continuously refined and the best path is selected. However, these approaches face significant challenges in expansive spaces, where increasing dimensionality causes the search space to grow substantially relative to regions of high information reward. In order to address these challenges, [30] introduced a sampling-based approach for IPP in large and high-dimensional search spaces. The method performs informed sampling within the continuous space and incorporates edge information gain during reward estimation, efficiently generating global paths that optimize information gathering within budget constraints.
@@
While [30] demonstrated effective global planning, the approach generated static plans that remained unchanged during robot execution. Our proposed algorithm, IA-TIGRIS, also uses a sampling-based framework but continuously optimizes and refines plans online as the robot acquires new information. This adaptive refinement is crucial for long-horizon planning scenarios where optimal paths are difficult to determine in a single planning cycle. Through improvements in tree-building logic and novel belief space node embeddings, IA-TIGRIS achieves sufficient computational efficiency for onboard deployment. We validate our method through extensive simulation testing and demonstrate efficient online planning capabilities on hardware across two distinct UAV systems.
@@ III Problem Formulation
This work addresses the challenge of maximizing data gathered by a robot within the specified budget constraints of the system, a problem formally known as the IPP problem. Our problem formulation is defined as
@@
We let $\mathcal{T}$ represent the sensor trajectory within the set of feasible trajectories $\mathbb{T}$, and $C(\mathcal{T})$ be the cost function for a given trajectory. An example of potential costs would be trajectory length, energy usage, or time and would not exceed the budget constraint $B$. Lastly, $I(\mathcal{T})$ is the reward or information gain from a trajectory, and an optimal trajectory $\mathcal{T}^{*}$ is found that maximizes the information gain and does not exceed the budget $B$.
@@ IV Proposed Approach
Figure 2: Overview of the IA-TIGRIS planning framework. As the robot is executing the current path, the belief map is updated using the current robot position. Whenever a new plan is initiated, the current robot position, belief map, and remaining budget are passed to the planner. The new path is merged into the previous path and given to the autopilot for execution.
This section details our proposed approach for informative path planning online using incremental and adaptive methods. A high level overview of our framework is shown in Fig. 2. Subsection IV-A provides a thorough explanation of the core algorithm, including the main planning loop and the graph update procedure, elucidating the steps involved in sampling, steering, and constructing the tree of potential paths. Subsection IV-B discusses the strategy to enable adaptive replanning with the presented algorithm. Finally, Subsection IV-C explores the techniques used to incrementally update the path, enhancing the efficiency and responsiveness of the planning process. Subsections IV-D and IV-E explain how the belief space and observation model integrate into the planner, and Subsection IV-F presents our reward models.
@@ IV-A Algorithm Overview
At the beginning of planning, the set of vertices $V$ is empty. The initial information $I_{s}$ at the start pose $x_{start}$ is calculated by $I(\cdot)$, and the initial cost $C$ is set to 0. The tuple $n_{start}$ is initialized and added to the set of vertices $V$, while the set of edges $E$ and the closed set of vertices $V_{closed}$ are initialized as empty. The graph G is initialized as $G\leftarrow(V,E)$.
The Algorithm 1 then loops until the computation time exceeds the allowed planning time $T$. During each iteration of the loop, a sample $x_{sample}$ is drawn from the state space $\mathcal{X}$ using the InformedSample function which is a weighted sampler based on the reward for viewing a state in the belief space. This is done by first conducting the weighted sample in the belief space and then uniformly sampling in the subset of the configuration space that views the sampled belief space state while respecting desired observation constraints. Informed sampling focuses paths on areas of higher reward rather than uniformly sampling in the search space.
@@ 11       UpdateGraph$(\mathbf{x}_{start},B)$
14while computation time $<T$ do
15       $\mathbf{x}_{sample}\leftarrow$ InformedSample$(\mathcal{X})$
@@ 17       $\mathbf{x}_{feas},e_{feas}\leftarrow$Steer$(n_{nearest},\mathbf{x}_{sample},\Delta,X_{free})$
18       AddPoseToGraph$(\mathbf{x}_{feas},e_{feas},n_{nearest},\Delta,R,B)$
@@ 23                  $\mathbf{x}_{new},e_{new}\leftarrow$Steer$(n_{near},\mathbf{x}_{feas},\!\Delta,X_{free})$
The nearest node $n_{nearest}$ in the graph $G$ to $x_{sample}$ is identified, and the algorithm attempts to steer from $n_{nearest}$ to $x_{sample}$ within a specified distance $\Delta$ while checking for collisions, resulting in a feasible pose $x_{feas}$ and an edge $e_{feas}$. In contrast with [30], the new pose and edge are immediately evaluated to be added to the graph $G$ using the AddPoseToGraph function, allowing the neighbor selection radius $R$ to be less than the extend distance $\Delta$.
The algorithm then searches for nearby nodes $N_{near}$ within a radius $R$ of $x_{feas}$ that are not in the closed set $V_{closed}$. For each neighbor $n_{near}$ in $N_{near}$, if $n_{near}$ is distinct from $x_{feas}$ and $n_{nearest}$, the algorithm attempts to steer from $n_{near}$ to $x_{feas}$. If the new pose $x_{new}$ is distinct from $n_{near}$, the pose and edge evaluated to be added to the graph $G$ using the AddPoseToGraph function. This neighborhood expansion structure is different than [30] in that it now allows for neighboring nodes to extend toward $x_{feas}$ even if it did not add and edge connecting to $n_{nearest}$.
After the computation time is exhausted, the algorithm extracts the best path $T$ from the graph $G$ using the BestPath function. This function simply returns the path with the highest information reward in the graph.
The AddPoseToGraph function, outlined in Algorithm 2, handles the addition of a new pose $x_{new}$ and edge $e_{new}$ to the graph $G$. The information $I_{new}$ for the new pose is calculated, and the cumulative cost $C_{new}$ is computed by adding the cost of the edge $e_{new}$ to the cost of $n_{parent}$. A new node $n_{new}$ is created as a tuple containing the new pose, information, and cumulative cost. The node $n_{new}$ is then checking by the Prune function. The pruning function checks nearby nodes within a radius to see if they contain a better solution. Because most of the nodes in the tree are not full paths, the upper bound on their rewards are compared. For more aggressive pruning, heuristics can be used for this comparison. In our case we use the heuristic of checking if a nearby node has a lower cost and higher current reward, but this heuristic can have implications on finding the optimal solution. If $n_{new}$ passes the pruning check, the edge $e_{new}$ and node $n_{new}$ are added to the sets $E$ and $V$, respectively, and graph $G$ is updated. If the cumulative cost $C_{new}$ equals the budget $B$, the new node $n_{new}$ is added to the closed set $V_{closed}$.

### v2

@@ Abstract
Planning paths that maximize information gain for robotic platforms has wide-ranging applications and significant potential impact. To effectively adapt to real-time data collection, informative path planning must be computed online and be responsive to new observations. In this work, we present IA-TIGRIS, an incremental and adaptive sampling-based informative path planner designed for real-time onboard execution. Our approach leverages past planning efforts through incremental refinement while continuously adapting to updated belief maps. We additionally present detailed implementation and optimization insights to facilitate real-world deployment, along with an array of reward functions tailored to specific missions and behaviors. Extensive simulation results demonstrate IA-TIGRIS generates higher-quality paths compared to baseline methods. We validate our planner on two distinct hardware platforms: a hexarotor UAV and a fixed-wing UAV, each having different motion models and configuration spaces. Our results show up to a 41% improvement in information gain compared to baseline methods, highlighting the planner’s potential for deployment in real-world applications. Project website and video at ia-tigris.github.io.
@@ I Introduction
Robots play a vital role in gathering information from the physical world, supporting a wide range of applications such as scientific research [1, 2, 3], environmental monitoring [4, 5, 6, 7, 8], search and rescue operations [9, 10], and disaster response efforts [11, 12, 13]. By employing intelligent algorithms, robotic systems enhance the efficiency of data collection, provide valuable insights, and support well-informed decision-making processes. These autonomous robots provide unparalleled advantages in situations where human access is constrained, dangerous, or logistically challenging [14]. Moreover, the scalability of autonomous information gathering robots enhances the rate of information gathering by not capping the number of robots based on the number of human operators.
@@
Informative path planning (IPP) seeks to create intelligent paths for robots that maximize their inherent potential and advantages in data gathering situations, while not violating budget constraints on the system or mission. This approach has many advantages over most coverage-based approaches which often don’t take into account the flight time of the drone or prior information about the space and don’t adapt to learned information during flight. However, solving the IPP problem is computationally expensive, being at least NP-hard even in a discrete space [15], leading to methods that only plan over a short horizon or are static during the path execution.
Recent works in informative path planning have explored learning-based approaches [16, 17] that aim to reason over longer planning horizons, as well as graph-based methods that optimize routes over fixed discretizations of the space [18, 19]. Although these approaches have advanced the field, many still rely on fixed-resolution graphs that can be difficult to scale to large, high-dimensional spaces. They also often use simplified sensor models (e.g., pointwise measurements or circular downward-facing fields of view), assume independent observation rewards without accounting for diminishing returns from repeated observations, and may not fully capture information gathered along continuous trajectories. In addition, learning-based methods can struggle when environmental conditions differ from those encountered during training [20].
To address these challenges, [21] presented an informed sampling-based algorithm that directly optimizes over the current belief map without requiring prior training. The approach scales to high-dimensional spaces, supports realistic sensor configurations, reasons over edge rewards to capture information gathered along the entire trajectory, and accounts for reward dependence, including multiple observations of the same state. However, plans were static and had to be rebuilt from scratch whenever the environment changed.
In this work, we introduce IA-TIGRIS (Incremental and Adaptive Tree-based Information Gathering using Informed Sampling), an adaptive and incremental sampling-based informative path planner. IA-TIGRIS builds on and substantially improves the approach in [21] by introducing new tree-building logic and novel belief map node embeddings, enabling efficient incremental planning as new information becomes available. IA-TIGRIS operates fully online, efficiently updating and building its set of possible trajectories as new information is gathered, directly incorporating changes in the belief map. This adaptive, incremental approach leads to global paths that are more robust to path execution disturbances, better adapt to evolving environment knowledge, and achieve higher quality than a single static plan through continual online refinement. Fig. 1 demonstrates IA-TIGRIS deployed on a hexarotor drone for an IPP application.
Our key contributions are the following:
@@
The paper is organized as follows: Section II gives a detailed overview of related works and the contributions of this work. Section III presents the problem formulation. Our proposed approach is detailed in Section IV. That section outlines the planner’s design and its integration with system components to enable adaptive and efficient informative path planning for real-world applications. Section V provides our detailed experimental evaluation and results of the approach. The results of our field deployments are summarized in Section VI, and Section VII offers the conclusion and future work.
@@ II Related Works
One way to formulate the IPP problem is to discretize the continuous space into a fixed set of states. Many previous works have then used this discrete space to view IPP as an orienteering problem. Given a weighted-undirected graph, orienteering involves determining a Hamiltonian path over a subset of nodes that maximizes total reward while satisfying a path cost constraint. The orienteering problem combines the objective of selecting a subset of nodes from the graph and solving a traveling salesman problem to minimize distance while visiting the selected nodes. The optimization objective for node selection can be tailored to specific applications. For example, [22] applies it to environmental monitoring in an aquatic settings while [23] applies it to the more general Constraint Satisfaction Problem (CSP).
@@
However, framing IPP as an orienteering problem leads to computationally expensive runtimes, which rapidly become intractable as the likelihood of near-optimal paths in the solution space diminishes. As [24] notes, the IPP problem is a NP-hard search problem. To overcome this, they develop an approximation algorithm that efficiently finds near-optimal solutions by leveraging mutual information in the problem space. They ensure that the mutual information formulation is submodular in nature [25]. This creates a diminishing returns property where each additional observation provides less new information than the previous one. Hence, the IPP problem can be reformulated as maximizing a submodular objective function while constrained to a budget. [26] studies this formulation by developing a recursive-greedy algorithm with strong theoretical guarantees.
@@
However, these approaches have only been used in applications that have relatively small search spaces such as environmental monitoring [22], water quality analysis [27], or chemical plume detection [28] that are discretized into a finite set of sensing locations. This becomes quickly intractable in larger, higher-dimensional applications such as searching for targets over large areas (e.g., missing hikers) or tracking multiple dynamic targets (e.g., ships at sea). To address this, [22] propose using heuristics while [22, 29, 26] suggest using greedy approaches to approximate solutions. Mixed Integer Programming (MIP) solutions [30] have also been proposed. But these solutions fail to capture the relationship between nodes which would be problematic for many applications. The limitations in capturing rewards relationships among nodes, combined with inefficient runtimes, necessitate developing better exploration algorithms for robust information gathering by autonomous systems in complex settings.
@@
Some previous approaches have used receding horizon-based methods to solve the IPP problem. Local receding-horizon solutions optimize the information objective over a small lookahead horizon, relying on replanning to approach globally optimality. One approach involves a receding horizon algorithm [31] that satisfies temporal logic specifications for safety-critical applications. Another approach [32] uses a receding horizon planner for informative planning in a latent environment modeled using Gaussian processes. According to [33], receding horizon-based path planners are susceptible to becoming trapped in local optima. To address this, they explore adjusting horizon length and propose a J-Horizon algorithm that incorporates lookahead step size to improve convergence toward better optima. Despite these improvements, receding horizon methods primarily plan optimal paths reactively and fail to fully utilize the available prior information [34].
@@
Notably, [35] compares lawnmower paths with other planning algorithms and found that lawnmower paths perform nearly as well as adaptive algorithms and outperform graph-based search algorithms that struggle to adapt to uncertainties in prior information. The ability to allow for revisits is very important for finding globally optimal paths, especially when the information objective is submodular or when the task requires revisits [21]. Branch and bound techniques show promise by pruning suboptimal branches early in the tree search [36, 37], but efficiently calculating tight bounds unknown, high-dimensional environments remains challenging. While Gaussian processes (GP) have been used to address this, implementing GPs is application specific and the computational complexity scale cubically with the number of observations [38]. Although methods exist to mitigate this complexity, long-horizon planning remains computationally demanding unless stationary assumptions are made about the covariance function. This is further compounded when planning in higher-dimensional state spaces.
Recent studies have explored learning-based and optimization-based approaches for enabling long-horizon informative path planning. Learning-based methods leverage offline-trained neural networks or policies to guide online decision-making. Specifically, [39] enhance online tree search by utilizing a network trained offline to efficiently sample informative robot actions. [40] develop a learning-based controller that dynamically switches among multiple classical local search strategies to enhance robustness in various scenarios. [16] presents CAtNIPP, an attention-based network trained to select subsequent robot states within a predefined sensing graph, while [41] extend this concept to dynamically select actions within an online-generated sensing graph. Optimization-based approaches explicitly model the information gain as an optimization objective and utilize numerical optimization techniques for planning. Examples include mixed-integer programming [19], Monte Carlo Tree Search (MCTS) [42], and dynamic programming [18]. Notably, the dynamic programming approach in [18] achieves efficiency through approximation strategies, such as assuming the robot teleports directly to future positions and calculating information gain based solely on current map entropy. However, most existing methods assume simplified point-based sensor models, significantly differing from realistic scenarios where sensor footprints are non-trivial, application-specific, and governed by practical sensor parameters. Moreover, these approaches generally overlook the robot’s heading and kinematic constraints, which substantially influence real-world planning outcomes.
@@
Numerous works address these drawbacks by formulating the task of information gathering in continuous space and introducing sampling-based algorithms. Examples of some of these works include [43], [44], [45, 46, 47, 48, 49, 50]. These methods employ sampling-based approaches, which involves the selection and inclusion of new states in continuous space, whether visiting all of the sampled states or adding them into a tree structure of potential trajectories. As the tree grows, the optimal path—maximizing information gain while respecting budget constraints—is continuously refined and the best path is selected. However, these approaches can face significant challenges in expansive spaces, where increasing dimensionality causes the search space to grow substantially relative to regions of high information reward. In order to address these challenges, [21] introduces a sampling-based approach for IPP in large and high-dimensional search spaces. The method performs informed sampling within the continuous space and incorporates edge information gain during reward estimation, efficiently generating global paths that optimize information gathering within budget constraints.
Another crucial aspect of IPP planners is their broad categorization into adaptive and non-adaptive planners. Following prior works such as [51, 20, 45, 52, 53, 41], we define adaptive planning methods as those that do not follow a static, predefined path, but instead continue to modify the plan during execution. This adaptivity enables the planner to respond to discrepancies between expected and observed measurements, to adjust for execution-time disturbances, and to refine the plan online as more information becomes available. Adaptivity is often implemented through frequent re-planning and belief updates. While this approach can improve performance, it increases computational complexity and requires lightweight planners capable of real-time execution on robotic platforms
@@
Our proposed algorithm, IA-TIGRIS, is an adaptive planner that builds upon the work of [21], which had demonstrated effective global planning, but generated static plans that remained unchanged during robot execution. IA-TIGRIS also uses a sampling-based framework but continuously optimizes and refines plans online as the robot acquires new information. This adaptive refinement is crucial for long-horizon planning scenarios where optimal paths are difficult to determine in a single planning cycle. In order to achieve sufficient computational efficiency for onboard deployment, we introduce improvements in the tree-building logic and add novel belief map node embeddings that facilitates efficient incremental planning. We validate our method through extensive simulation testing and demonstrate efficient online planning capabilities on hardware across two distinct UAV systems.
@@ III Problem Formulation
This work addresses the challenge of maximizing data gathered by a robot within the specified budget constraints of the system, a problem formally known as the informative path planning (IPP) problem. Following [21], we define it as
@@
Here, $\mathcal{T}$ represent the sensor trajectory from the set of feasible trajectories $\mathbb{T}$, and $C(\mathcal{T})$ is its cost. Example of potential cost functions include trajectory length, energy consumption, or flight time, and each would not exceed the budget constraint $B$. Lastly, $I(\mathcal{T})$ denotes the reward or information gain from trajectory $\mathcal{T}$, and the goal is to find the optimal trajectory $\mathcal{T}^{*}$ that maximizes information gain without exceeding the budget.
In this work, trajectories are planned in a four-dimensional space defined by $(x,y,z,\psi)$, where $x,y,z$ specify the robot’s position and $\psi$ its heading. We typically sample new states from a discrete set of altitudes (for $z$), while $x,y,\psi$ remain continuous. In our experiments, the cost $C(\mathcal{T})$ is the total distance traveled by the robot. The information gain function $I(\mathcal{T})$ computes expected reduction in uncertainty over the environment, modeled as a belief map representing the probability of object presence across spatial cells. While in our applications this belief map encodes discrete objects of interest, the algorithm itself is agnostic to the underlying representation. Other implementations could use different belief models, environment features, or custom metrics for information gain. Importantly, our reward function is not constrained to be modular or submodular, allowing for more expressive and task-specific definitions of information gain.
@@ IV Proposed Approach
Figure 2: The IA-TIGRIS planning framework consists of two main components: the belief map and the informative path planner. The belief map is responsible for managing the agents understanding of the environment while the informative path planner generates adaptive plans based on the agents position and remaining budget. For the planning procedure, the new path is merged with the existing plan by pruning outdated portions of the existing path and starting the new plan from the agents expected position.
This section details our proposed approach for informative path planning online using incremental and adaptive methods. Subsection IV-A provides a thorough explanation of the core algorithm, including the main planning loop and the graph update procedure, elucidating the steps involved in sampling, steering, and constructing the tree of potential paths. Subsection IV-B discusses the strategy to enable adaptive replanning with the presented algorithm. Finally, Subsection IV-C explores the techniques used to incrementally update the path, enhancing the efficiency and responsiveness of the planning process. Subsections IV-D and IV-E explain how the belief map and observation model integrate into the planner, and Subsection IV-F presents our reward models. A high level overview of this framework is shown in Fig. 2.
@@ 7    $V\leftarrow{n_{start}}$; $E\leftarrow\emptyset$; $V_{closed}\leftarrow\emptyset$; $G\leftarrow(V,E)$
14while computation time <T<T do
[…]
@@ 17    $\mathbf{x}_{feas},e_{feas}\leftarrowSteer$$(n_{nearest},\mathbf{x}_{sample},\Delta,X_{free})$
[…]